OpenStudy (anonymous):

what is the domain of f(x)=cos(x)? what is the y value of tan(x) when x=-180 the range of f(x)=sin(x)is−1≤y≤1 soo sorry about all the questions I really don't understand this section

OpenStudy (anonymous):

@SomeGirl1999 help?

OpenStudy (anonymous):

@Thatyoungbuttercup help?

OpenStudy (anonymous):

For the first question the answer is (-infinity,+infinity)

OpenStudy (somegirl1999):

Sorry I'm not good at math.

OpenStudy (anonymous):

the choices are a.-1 b.0 c.1 d.does not exist

OpenStudy (anonymous):

@SomeGirl1999 ok thanks anyways

OpenStudy (anonymous):

oh wait no those answers are for a different quiestion

OpenStudy (anonymous):

??? so the answer choices are? lol

OpenStudy (anonymous):

a.all real numbers b.\[-1\le x \le1\] c.\[x \le 0\] d.\[x \ge 0\]

OpenStudy (anonymous):

for the first question?

OpenStudy (anonymous):

yeah

OpenStudy (shubhamsrg):

I dont think you should be told the direct answer,

OpenStudy (shubhamsrg):

because that is simply cheating :|

OpenStudy (anonymous):

^True! but i dont know how to explain it though:(

OpenStudy (shubhamsrg):

nothing to explain in the domain of cosx. @nikakuz should better pick up a book and learn as a fact, the domains of basic trigonometric fuctions.

OpenStudy (shubhamsrg):

functions*

OpenStudy (anonymous):

I did I tried but I really don't understand it Im horrible at math

OpenStudy (shubhamsrg):

cheating is not gonna help

OpenStudy (shubhamsrg):

so, what do you understand by the term domain ?

OpenStudy (anonymous):

like x?

OpenStudy (shubhamsrg):

hmm, close enough. It means the values x can take such that it gives out a real value of y. in your case, y=cosx Do you think there is any value of x, on the number line, for which there is no real value of y ?

OpenStudy (anonymous):

um no?

OpenStudy (shubhamsrg):

hmm right, so your answer will be... ?

OpenStudy (anonymous):

i don't know i don't get how to get the answer from that

OpenStudy (shubhamsrg):

yes , you don't. And that is just sad . :(

OpenStudy (anonymous):

yeah cuz im trying to learn this from an online class so its hard and im terrible at math so that doesn't help

OpenStudy (anonymous):

please just help me i did a bunch of other problems i just don't understand these 3

OpenStudy (anonymous):

@Thatyoungbuttercup please help me

OpenStudy (anonymous):

@shubhamsrg dude help a boy out! he has been working hard just cause he doesnt show it doesnt mean he has @nikakuz i forgot all this sorry i just know the answer for the first one but i dont know how to explain

OpenStudy (anonymous):

the first question is all real numbers!

OpenStudy (anonymous):

alright thank you!!

OpenStudy (anonymous):

wait i think i got the answer for the second one

OpenStudy (anonymous):

the choices i gave earlier are for the second one

OpenStudy (anonymous):

like the first choices

OpenStudy (anonymous):

what!0_o what is the domain of f(x)=cos(x)? ANSWER:all real numbers! what is the y value of tan(x) when x=-180 ANSWER:0

OpenStudy (anonymous):

nvm haha thank you soo much

OpenStudy (shubhamsrg):

Well you can;t learn unless you make effort from your side too. You are not making effort to learn. hmm, nevermind, I'd still like to wish you good luck for the future.

OpenStudy (anonymous):

@nikakuz You got it now kinda for the last one?

OpenStudy (anonymous):

@shubhamsrg i tried making an effort but i just don't understand it

OpenStudy (anonymous):

not really true or false though so at least i have a 50/50 chance

OpenStudy (anonymous):

its true!

OpenStudy (anonymous):

yayy thank you soo much i just submitted the assignment and i got 20 out of 20 yayy

OpenStudy (anonymous):

Im so happy for you dude! :D

OpenStudy (anonymous):

:D